210 J of heat is added to a system having an internal energy of 58 J. Calculate the amount of the external work done by the system.


1
Expert's answer
2022-01-19T06:00:02-0500

From the first law of thermodynamics

"\Delta U=Q+W"

"\Delta U=58J"

"Q=210J"

"W=?"

"\therefore W=58-210 = -152J"

The amount of external work done is 152J